the sum of two consecutive even integers is 94
the sum of three consecutive odd integers is 123

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

n+n+2=94
2n=92
n=46
n+2=48
n+n+2+n+4=123
3n+6=123
3n=117
n=39
n+2=41
n+4=43

Answer by Alan3354(69443)

And? Did you want to find the integers?
------------------
the sum of two consecutive even integers is 94
94/2 = 47
---> 46 & 48
-----------------
the sum of three consecutive odd integers is 123
123/3 = 41
--> 39, 41, 43